# Turning to Online Forums for Legal Information: A Case Study of GDPR’s Legitimate Interests

### v1

@@ 6 Discussion
Posters are motivated to comply but still need guidance in applying legitimate interests in practical settings. It does not seem like most posters are actively trying to circumvent the law based on the amount of posts seeking legal information, and the amount of views posts had. Rather, our analysis suggests there is a wide range of potential areas of confusion and doubt amongst practitioners, therefore showing that practitioners require more guidance when applying legitimate interests and the GDPR in practice.
@@
Time and investment from commentators might give credibility. In our analysis, most Accepted answers were considered as legally sound based on the information provided. Comprehensive answers (with case law notes, excerpts of guidelines, and references to national laws) denote time and attention from commentators. Unlike Stack Overflow, which can be used in job applications to display one’s technical knowledge and abilities [80], those answering Law Stack Exchange questions did not generally specify if they had a legal background, nor it was clear what is the incentive for commentators to participate and give thorough answers for free.
Dedicated resources exist but are not referred to in forums. Online forums may be preferred by practitioners to find legal information because of the granularity of questions and answers which relate to applying the GDPR for a specific scenario. The legal sources that were cited in the answers do not mention already-available online resources for developers mentioned in Section 2.2.1, but rather to DPA guidelines. The granularity of concrete scenarios that developers face seem to have inspired the creation of the CNIL’s developer code, which is an example of how the law can be operationalized to more concrete settings. Interestingly, this resource was not mentioned in posts or comments we analyzed.
@@ 6.1 Implications for Legal Compliance
Resorting to these forums may indicate a lack of legal resources for smaller and medium-sized companies who in turn might not have in-house legal teams or department at their disposal to discuss GDPR compliance issues. We posit that the status and the social responsibility of online legal forums merit attention, and call for discussions on the features that should be added or removed to make the sourced information more helpful, accurate, complete, and shaped to the practitioners’ needs.
@@ 6.2 Guidelines Towards Writing Forum Posts for Legally-Sound Answers
Therefore, it is important that online forums provide correct legal information for current and future uses of this information for legal compliance. Our analysis found that most commentators take time and care in crafting their responses, but legal soundness is also reliant on other factors, such as a poster’s jurisdiction, practices and industry. In this section, we present a three recommendations to ensure that posters can receive responses that are as legally sound as possible, which online legal forums could include in their rules and guidelines for posters.
First, posters should include the jurisdiction of their organization, and the jurisdiction of their user base. The jurisdiction should specify the country at stake, since different European Member States may have different national law specifications. Second, posters should also mention i) the industry their organization is involved in, especially if it deals with sensitive data, such as health information, so that posters know of special safeguards they need to adopt with users’ data, ii) processing purposes, and iii) their concrete role within data processing activities. Third, commentators should provide a disclaimer about the limitations of the concrete answer and add the fact that the dedicated legal information should neither bound the poster nor be construed or interpreted as legal advice.
@@ 6.3.1 Future directions
As GDPR compliance involved several stakeholders not only limited to those in law, but also those in tech, we suggest more collaborations between HCI researchers and the legal community to better understand the human factors of legal compliance and create solutions for practitioners involved.
@@ 7 Conclusion
We believe that, while there exists some organizations with malicious intent for using the legitimate interest legal basis to collect user data, legitimate interests are prone to being misused due to the broad and flexible nature, leading practitioners to misunderstand its concrete applications. In light of this, the CHI and legal communities should consider the role of online forums in the data protection landscape, and pay more attention to the human factors of legal compliance to give more actionable and scenario-specific guidelines to foster data protection by design and by default.

### v2

@@ 6 Discussion
Posters are motivated to comply but still need guidance in applying legitimate interests in practical settings. Those who do not care about compliance or are malicious in their compliance are unlikely to seek legal information in the first place. However, of those who did post a question, it does not seem like most posters are actively trying to circumvent the law based on the amount of posts seeking legal information, and the amount of views posts had. Rather, our analysis suggests there is a wide range of potential areas of confusion and doubt amongst practitioners, therefore showing that practitioners require more guidance when applying legitimate interests and the GDPR in practice.
@@
Time and investment from commentators might give credibility. In our analysis, most Accepted answers were considered as legally sound based on the information provided. Comprehensive answers (with case law notes, excerpts of guidelines, and references to national laws) denote time and attention from commentators. Unlike Stack Overflow, which can be used in job applications to display one’s technical knowledge and abilities [91], those answering Law Stack Exchange questions did not generally specify if they had a legal background, nor was it clear what the incentive is for commentators to participate and give thorough answers for free.
Dedicated resources exist but are not referred to in forums. Online forums may be preferred by practitioners to find legal information because of the granularity of questions and answers which relate to applying the GDPR for a specific scenario. The legal sources that were cited in the answers do not mention already-available online resources for developers mentioned in Section 2.2.1, but instead referred to DPA guidelines. The granularity of concrete scenarios that developers face seem to have inspired the creation of the French Data Protection Authority’s (CNIL) developer guide [23], which is an example of how the law can be operationalized to more concrete settings. Interestingly, this resource was not mentioned in posts or comments we analyzed.
@@ 6.1 Implications for Legal Compliance
Resorting to these forums may indicate a lack of legal resources for smaller and medium-sized companies who in turn might not have in-house legal department at their disposal to discuss GDPR compliance issues. We posit that the status and the social responsibility of online legal forums merit attention, and call for discussions on the features that should be added or removed to make the sourced information more helpful, accurate, complete, and shaped to the practitioners’ needs.
@@ 6.2 Guidelines Towards Writing Forum Posts for Legally-Sound Answers
Therefore, it is important that online forums provide correct legal information for current and future uses of this information for legal compliance. Our analysis found that most commentators take time and care in crafting their responses, but legal soundness is also reliant on other factors, such as a poster’s jurisdiction, practices and industry. In this section, we present three recommendations to ensure that posters can receive responses that are as legally sound as possible, which online legal forums could include in their rules and guidelines for posters.
First, posters should include the jurisdiction of their organization, and the jurisdiction of their user base. The jurisdiction should specify the country at stake, since different European Member States may have different national law specifications. Second, posters should also mention i) the industry their organization is involved in, especially if it deals with sensitive data, such as health information, so that posters know of special safeguards they need to adopt with users’ data, ii) processing purposes, and iii) their concrete role within data processing activities. Third, commentators should provide a disclaimer about the limitations of the concrete answer and add the fact that the dedicated legal information should neither bind the poster nor be construed or interpreted as legal advice.
@@ 6.3.1 Future directions
As GDPR compliance involved several stakeholders not only limited to those in law, but also those in tech, we suggest more collaborations between human-computer interaction researchers and the legal community to better understand the human factors of legal compliance and create solutions for practitioners involved.
@@ 7 Conclusion
We believe that legitimate interests are prone to being misused due to the broad and flexible nature, leading practitioners to misunderstand their concrete applications. In light of this, the human-computer interaction and legal communities should consider the role of online forums in the data protection landscape, and pay more attention to the human factors of legal compliance to give more actionable and scenario-specific guidelines to foster data protection by design and by default.